OpenStudy (anonymous):

I know the answer for; -12x-20=10x+50 is x=-35, but can someone please explain the steps to getting this answer?

hartnn (hartnn):

add 12x to both sides, what u get ?

hartnn (hartnn):

-20=10x+12x+50 can u solve further?

OpenStudy (anonymous):

for the first example, I get -4 on one side and 82 for the second side:(

hartnn (hartnn):

82? i'll write the steps, see whether u get it : -12x-20=10x+50 Add 12x to both sides -20 = 10x +12 x +50 -20 = 22 x + 50 subtract 50 from both sides, -20-50 = 22x 22x = -70 x = -70/ 22 = -35 /11 is it +10x or -10x on right??

hartnn (hartnn):

if it is -10 x only then u get x=-35

OpenStudy (anonymous):

Thank you very much - I couldn't find this in my book at all . Now I can work numerous examples and finally understand the process!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

ask if there is any doubt in any step.